Originally the dimensions of a rectangle or 20 cm x 23 cm when the dimensions were decreased of the same amount the area of the rectangle decrease by 120 cm² find the dimensions of the new rectangle
Answer by mananth(16949)

the original dimensions 20 * 23
let the diminished length be x
new dimensions (23-x) & (20-x)
New Area =20*23-120
=340 cm^2
(20-x)(23-x)=340
460-20x-23x+x^2=340
x^2-43x+120=0
x^2-40x-3x+120=0
x(x-40)-3(x-40)=0
(x-40)(x-3)=0
x=3 OR 40
40 is not possible
the width reduced is 3 cm
New dimension 17 by 20 cm